OpenStudy (rianiscool):

need help asap in math Use properties to evaluate the expression. (2 • 9) • 4 A. 72 B. 54 C. 38 D. 15

OpenStudy (rianiscool):

will give a medal

OpenStudy (igreen):

\((2 \cdot 9) \cdot 4\) \((18) \cdot 4\) \(72\)

OpenStudy (igreen):

Remember, Parenthesis first.